OpenStudy (erinweeks):

Find the unit vector that has the same direction as the vector v. v = 3i + j

OpenStudy (erinweeks):

@AkashdeepDeb or @Psymon can you help?? (:

OpenStudy (psymon):

Yep. So just like we did in the last problem, we need the magnitude first. Which means we do m^2 = i^2 + j^2. Once you get the magnitude, divide both i and j by that magnitude. So your final answer looks like this: \[\frac{ vector (v) }{ ||v|| }\]

OpenStudy (akashdeepdeb):

Exactly! Because Magnitude . Unit vector = Actual vector! So unit Vector = Actual Vector/Magnitude !! :D

OpenStudy (psymon):

I need to brb, though. So if you got any questions just bug @AkashdeepDeb Ill be back xD

OpenStudy (akashdeepdeb):

-.-

OpenStudy (akashdeepdeb):

@ErinWeeks Did you get it though?

OpenStudy (akashdeepdeb):

@ErinWeeks ?

OpenStudy (akashdeepdeb):

?

OpenStudy (erinweeks):

sorry i was playing call of duty :0 lol not really because i dont get what to plug in

OpenStudy (akashdeepdeb):

What is the magnitude of the vector?

OpenStudy (akashdeepdeb):

@ErinWeeks ?

OpenStudy (erinweeks):

uhm im honestly not sure im looking in my book trying to see

OpenStudy (erinweeks):

wouldnt it be 3?

OpenStudy (akashdeepdeb):

It'd be sqrt 10 @Psymon 'd explain :D

OpenStudy (psymon):

Sorry if I wasnt clear before about it. The magnitude is the length of the vector. And that length is found using: \[(magnitude)^{2} = i ^{2}+j ^{2} \]. Its the same thing as pythagorean theorem, a^2 + b^2 = c^2, its just different letters. So to get magnitude, lets check. i is 3, j is 1 \[m ^{2}=3^{2}+1^{2}\] \[m ^{2}=9 + 1\] \[m ^{2}=10 -> m = \sqrt{10}\] That kind of make sense?

OpenStudy (psymon):

@AkashdeepDeb I wasnt out to just volunteer you earlier, you were already here and I had to brb xD

OpenStudy (akashdeepdeb):

Yeah I get it! But then he/she started playing call of duty or some game. XD

OpenStudy (psymon):

Magnitude of the stray bullet shot into the enemy player?

OpenStudy (erinweeks):

yea i makes more sense but now thats not the final answer .. & sorrry i was playing the new dlc maps :O

OpenStudy (psymon):

Right, thats just the magnitude. To get a unit vector, we divide i and j by the magnitude. \[\frac{ i }{ magnitude }+\frac{ j }{ magnitude }\] \[= \frac{ 3i }{ \sqrt{10} }+\frac{ j }{ \sqrt{10} } \] And that would be the unit vector we want.

OpenStudy (erinweeks):

oh okay i see i got it..

OpenStudy (psymon):

mhm ^_^ Glad it makes sense.

OpenStudy (erinweeks):

it actually makes alot more sense

OpenStudy (psymon):

Oh, and theres a way to check and see if its a unit vector, too :3